# Multipoint distributions of the KPZ fixed point with compactly supported initial conditions

### v2

@@ Abstract
In this paper, we obtain the first formula for the space-time joint distributions of the KPZ fixed point with general initial conditions of compact support. The formula is obtained through taking $1:2:3$ KPZ scaling limit of the multipoint distribution formulas for the totally asymmetric simple exclusion process (TASEP). A key novelty is a probabilistic representation, inspired by [MQR21], of the kernel encoding the initial condition for TASEP, which was first defined through an implicit characterization in [Liu22]. Moreover, we also verify that the equal time degenerated version of our formula matches the path integral formula in [MQR21] for the KPZ fixed point.
@@ 1.1 Background
It is well known (see, e.g., [BDJ99, Joh00, ACQ11]) that for special initial conditions, the one point marginals of $\mathcal{H}(\alpha,\tau)$ are described by the Tracy-Widom distribution and its relatives. Extensions to joint laws of multiple spatial points at equal time was obtained in [PS02, BFPS07, BFS08, BFP10], leading to explicit descriptions of the spatial process $\mathcal{H}(\cdot,\tau)$ for special initial conditions. In the breakthrough work [MQR21], the authors were able to find explicit Fredholm determinant formulas for the joint laws of $\mathcal{H}(\alpha_{1},\tau;\mathfrak{h}),\cdots,\mathcal{H}(\alpha_{m},\tau;\mathfrak{h})$, starting from general upper semicontinuous initial conditions. This leads a complete description of the Markovian dynamics of the fixed point.
@@ 1.2 Main results
For technical reasons, We will mostly work with the dense subspace of $\mathrm{UC}$ consisting of functions that are -$\infty$ outside of a compact set.
@@
Our main results are formulas for the joint distributions of the KPZ fixed point starting with initial condition $\mathfrak{h}\in\mathrm{UC}_{c}$, at arbitrary many distinct space-time points $(\alpha_{1},\tau_{1}),\cdots,(\alpha_{m},\tau_{m})$. To state the result, we introduce the following total ordering $\prec$ on the space-time plane $\mathbb{R}\times\mathbb{R}$:
@@
Let $\mathfrak{h}\in\mathrm{UC}_{c}$. Then for any $m\geq 1$ and any $m$ space-time points $(\alpha_{1},\tau_{1})\prec\cdots\prec(\alpha_{m},\tau_{m})\in\mathbb{R}\times\mathbb{R}_{+}$, we have the following formula for the multi-point distribution of the KPZ fixed point $\mathcal{H}(\alpha,\tau;\mathfrak{h})$:
$\mathbb{P}\left(\bigcap_{\ell=1}^{m}\left\{\mathcal{H}(\alpha_{\ell},\tau_{\ell};\mathfrak{h})\leq\beta_{\ell}\right\}\right)=\oint_{0}\frac{\mathop{}\!\mathrm{d}z_{1}}{2\pi\mathrm{i}z_{1}(1-z_{1})}\cdots\oint_{0}\frac{\mathop{}\!\mathrm{d}z_{m-1}}{2\pi\mathrm{i}z_{m-1}(1-z_{m-1})}\mathbf{D}_{\mathfrak{h}}(z_{1},\cdots,z_{m-1}).$ (1.6)
The function $\mathbf{D}_{\mathfrak{h}}(z_{1},\cdots,z_{m-1})$ is defined as a Fredholm determinant in Definition 2.1. An equivalent definition through a series expansion will be discussed in Section 2.2.
@@
The function $\chi_{\mathfrak{h}}(\eta,\xi)$ is defined in terms of Brownian motion hitting expectations, an idea highly inspired by [MQR21]. Indeed, $\chi_{\mathfrak{h}}(\eta,\xi)$ should be understood as the Brownian hitting operators in [MQR21] written in Fourier-like spaces. Nevertheless we stress that our results do not follow directly from [MQR21]. In the multi-time situation, direct connections to determinantal point processes and the Eynard-Mehta theorem are lost and the bi-orthoganalization procedure here arises in a different way and takes a different form. On the contrary, our results are, in some sense, more general. Indeed if we set the time parameters to be the same (which is allowed in the assumption of the theorem), the right-hand side of (1.6) can be shown to recover the formulas in [MQR21], after some quite non-trivial manipulations. We refer to Section 6.2 for the details, see also [LO25] which treats the special narrow wedge case.
@@ 1.3 Outline of the proof and some discussions
A key novelty of this paper is that we find an explicit probabilistic representation of the function $\mathrm{ch}_{Y}(v,u)$, through a hitting expectation with respect to geometric random walks, see Theorem 3.4. The probabilistic representation is suitable for asymptotic analysis and leads to the Brownian hitting representation in the limit. For technical reasons, we first take the limit of the TASEP formula under the assumption that the KPZ fixed point starts with initial conditions consisting of finitely many narrow wedges, and then extend the formula to compactly supported initial condition at the level of the KPZ fixed point, using a density argument and the continuity of the law of the KPZ fixed point with respect to initial conditions.
Finally we comment on our assumptions on the initial conditions. It would be desirable if one can get a formula that works for all initial conditions $\mathfrak{h}\in\mathrm{UC}$, in particular, the flat initial condition $\mathfrak{h}\equiv 0$. The reason we choose to restrict to the subspace $\mathrm{UC}_{c}$ is not merely a technical issue. There are genuine structural difficulties in this generality: the characteristic function $\chi_{\mathfrak{h}}(\eta,\xi)$ of the initial condition (see Definition 2.2) may not be well-defined pointwisely in general. Indeed for the flat initial condition $\mathfrak{h}\equiv 0$, one can show from our formula that $\chi_{\mathfrak{h}}(\eta,\xi)$ is the dirac delta kernel $\delta_{\eta=-\xi}$, when properly interpreted. This suggests that in general, one should conjugate our formula to real spaces so that the limit when the support goes to infinity exists, at the level of operators acting on real spaces. We choose to stick to contour integral/Fourier type kernels as it makes the algebraic structure of the multi-time formula much nicer and the analogy between narrow wedge and general initial conditions much more transparent. We leave it as a future project to extend our formula to any $\mathfrak{h}\in\mathrm{UC}$, with a kernel acting on real spaces.
@@
Throughout the paper, we will mostly use english letters $x,t,h,u,v,w,\cdots$ for the pre-limit (TASEP) formulas and greek letters $\alpha,\tau,\beta,\xi,\eta,\zeta,\cdots$ for the limiting (KPZ fixed point) formulas. A detailed summary of the notation we use is in the following table.
@@
The rest of the paper is organized as follows. In Section 2 we present the formulas for the main part $\mathbf{D}_{\mathfrak{h}}$ appearing in Theorem 1.2, both as a Fredholm determinant in Section 2.1, and as a Fredholm series expansion in Section 2.2. Then in Section 3 we present and prove the corresponding pre-limit formulas for TASEP, in particular in Section 3.1 we prove that the characteristic function of the initial condition is given by a random walk hitting expectation. In Section 4 we prove convergence of the TASEP formulas to the KPZ fixed point formulas, under the assumption that the initial condition of the KPZ fixed point consists of finitely many narrow wedges. We then extend the KPZ fixed point formula to any compactly supported initial condition in Section 5. Finally in Section 6 we show that at equal-time, our formula reduces to a genuine Fredholm determinant, which is then shown to be equivalent to the path integral formula of [MQR21].
@@ Acknowledgements
YL acknowledges the support of a Junior Fellowship at Institut Mittag-Leffler in Djursholm, Sweden, during the Fall semester of 2024, funded by the Swedish Research Council under grant no. 2021-06594. The work of ZL is partially supported by NSF DMS-2246683. The authors thank Li-Cheng Tsai and Xuan Wu for their comments and suggestions.
@@ 2 Multipoint distribution formula for the KPZ fixed point
In this section we explain in details the function $\mathbf{D}_{\mathfrak{h}}(z_{1},\cdots,z_{m-1})$ appearing on the right-hand side of (1.6). Proofs of the formula will be deferred to Section 4 and Section 5.
2.1 Fredholm determinant representation of $\mathbf{D}_{\mathfrak{h}}(z_{1},\cdots,z_{m-1})$
The function $\mathbf{D}_{\mathfrak{h}}$ is defined in the same way as in [Liu22] for the narrow wedge initial condition, except for the part that involves the initial condition $\mathfrak{h}$ which only appears in the top-left corner of the integral kernel. Below we introduce the Fredholm determinant representation of the function $\mathbf{D}_{\mathfrak{h}}$.
@@ 2.1 Fredholm determinant representation of $\mathbf{D}_{\mathfrak{h}}(z_{1},\cdots,z_{m-1})$
Let $\Gamma_{m,\mathrm{L}}^{\mathrm{out}}$, $\cdots$, $\Gamma_{2,\mathrm{L}}^{\mathrm{out}}$, $\Gamma_{1,\mathrm{L}}$, $\Gamma_{2,\mathrm{L}}^{\mathrm{in}}$, $\cdots$, $\Gamma_{m,\mathrm{L}}^{\mathrm{in}}$ be 2m12m-1 “nested” contours in the region $\mathbb{C}_{\mathrm{L}}$. They are all unbounded contours from $\infty\mathrm{e}^{-2\pi\mathrm{i}/3}$ to $\infty\mathrm{e}^{2\pi\mathrm{i}/3}$. Moreover, they are located from the right (corresponding to the superscript $\mathrm{out}$”) to the left ($\mathrm{in}$”). The superscripts $\mathrm{out}$” and $\mathrm{in}$” should be understood with respect to the point -$\infty$. Similarly, let $\Gamma_{m,\mathrm{R}}^{\mathrm{out}}$, $\cdots$, $\Gamma_{2,\mathrm{R}}^{\mathrm{out}}$, $\Gamma_{1,\mathrm{R}}$, $\Gamma_{2,\mathrm{R}}^{\mathrm{in}}$, $\cdots$, $\Gamma_{m,\mathrm{R}}^{\mathrm{in}}$ be 2m12m-1 “nested” contours from left to right on the half plane $\mathbb{C}_{\mathrm{R}}$. They are from $\infty\mathrm{e}^{-\pi\mathrm{i}/5}$ to $\infty\mathrm{e}^{\pi\mathrm{i}/5}$. Their superscripts $\mathrm{out}$” and $\mathrm{in}$” could be understood with respect to the point $\infty$. Note that the angles for the left contours and right contours are chosen differently. The choice of the angles guarantees super-exponential decay of the kernel along the contours even if $\tau_{i}=\tau_{i+1}$ for some $i$. See [Liu22, LZ25] for more discussions on the choices of the angles. See Figure 1 for an illustration of the contours.
@@
$\Gamma_{\ell,\mathrm{L}}:=\Gamma_{\ell,\mathrm{L}}^{\mathrm{out}}\cup\Gamma_{\ell,\mathrm{L}}^{\mathrm{in}},\qquad\Gamma_{\ell,\mathrm{R}}:=\Gamma_{\ell,\mathrm{R}}^{\mathrm{out}}\cup\Gamma_{\ell,\mathrm{R}}^{\mathrm{in}},\qquad\ell=2,\cdots,m,$
@@
${\mathrm{d}}\mu(\zeta)={\mathrm{d}}\mu_{\bm{z}}(\zeta):=\begin{dcases}\frac{-z_{\ell-1}}{1-z_{\ell-1}}\frac{{\mathrm{d}}\zeta}{2\pi{\mathrm{i}}},&\zeta\in\Gamma_{\ell,\mathrm{L}}^{\mathrm{out}}\cup\Gamma_{\ell,\mathrm{R}}^{\mathrm{out}},\quad\ell=2,\cdots,m,\\ \frac{1}{1-z_{\ell-1}}\frac{{\mathrm{d}}\zeta}{2\pi{\mathrm{i}}},&\zeta\in\Gamma_{\ell,\mathrm{L}}^{\mathrm{in}}\cup\Gamma_{\ell,\mathrm{R}}^{\mathrm{in}},\quad\ell=2,\cdots,m,\\ \frac{{\mathrm{d}}\zeta}{2\pi{\mathrm{i}}},&\zeta\in\Gamma_{1,\mathrm{L}}\cup\Gamma_{1,\mathrm{R}}.\end{dcases}$ (2.2)
@@
$\mathrm{D}_{\mathfrak{h}}(z_{1},\cdots,z_{m-1})=\det\left(\mathrm{I}-\mathrm{K}_{1}\mathrm{K}_{\mathfrak{h}}\right),$
@@
$\mathrm{f}_{i}(\zeta):=\begin{dcases}\mathrm{e}^{-\frac{1}{3}(\tau_{i}-\tau_{i-1})\zeta^{3}+(\alpha_{i}-\alpha_{i-1})\zeta^{2}+(\beta_{i}-\beta_{i-1})\zeta},&i=2,\cdots,m,\\ \mathrm{e}^{-\frac{1}{3}\tau_{1}\zeta^{3}+\alpha_{1}\zeta^{2}+\beta_{1}\zeta},&i=1.\end{dcases}$ (2.7)

### v3

@@ Abstract
In this paper, we obtain the first formula for the space-time joint distributions of the KPZ fixed point with general initial conditions of compact support. The formula is obtained through taking $1:2:3$ KPZ scaling limit of the multipoint distribution formulas for the totally asymmetric simple exclusion process (TASEP). A key ingredient is a probabilistic representation, inspired by [MQR21], of the kernel encoding the initial condition for TASEP, which was first defined through an implicit characterization in [Liu22]. Moreover, we also verify that the equal time version of our formula matches the path integral formula in [MQR21] for the KPZ fixed point.
@@ 1.1 Background
It is well known (see, e.g., [BDJ99, Joh00, ACQ11]) that for special initial conditions, the one point marginals of $\mathcal{H}(\alpha,\tau)$ are described by the Tracy-Widom distribution and its relatives. Extensions to joint laws of multiple spatial points at equal time were obtained in [PS02, BFPS07, BFS08, BFP10], leading to explicit descriptions of the spatial process $\mathcal{H}(\cdot,\tau)$ for special initial conditions. In the breakthrough work [MQR21], the authors were able to find explicit Fredholm determinant formulas for the joint laws of $\mathcal{H}(\alpha_{1},\tau;\mathfrak{h}),\ldots,\mathcal{H}(\alpha_{m},\tau;\mathfrak{h})$, starting from general upper semicontinuous initial conditions. This leads to a complete description of the Markovian dynamics of the fixed point. We also remark that the results of [MQR21] were further generalized in [MR23a, BLSZ23, MR23b].
@@ 1.2 Main results
For technical reasons, We will mostly work with the dense subspace of $\mathrm{UC}$ consisting of functions that are -$\infty$ outside a compact set.
@@
Our main results are formulas for the joint distributions of the KPZ fixed point starting with initial condition $\mathfrak{h}\in\mathrm{UC}_{c}$, at arbitrary many distinct space-time points $(\alpha_{1},\tau_{1}),\ldots,(\alpha_{m},\tau_{m})$. To state the result, we introduce the following total ordering $\prec$ on the space-time plane $\mathbb{R}\times\mathbb{R}$:
@@
Let $\mathfrak{h}\in\mathrm{UC}_{c}$. Then for any $m\geq 1$ and any $m$ space-time points $(\alpha_{1},\tau_{1})\prec\cdots\prec(\alpha_{m},\tau_{m})\in\mathbb{R}\times\mathbb{R}_{+}$, we have the following formula for the multipoint distribution of the KPZ fixed point $\mathcal{H}(\alpha,\tau;\mathfrak{h})$:
$\mathbb{P}\left(\bigcap_{\ell=1}^{m}\left\{\mathcal{H}(\alpha_{\ell},\tau_{\ell};\mathfrak{h})\leq\beta_{\ell}\right\}\right)=\oint_{0}\frac{\mathop{}\!\mathrm{d}z_{1}}{2\pi\mathrm{i}z_{1}(1-z_{1})}\cdots\oint_{0}\frac{\mathop{}\!\mathrm{d}z_{m-1}}{2\pi\mathrm{i}z_{m-1}(1-z_{m-1})}\mathbf{D}_{\mathfrak{h}}(z_{1},\ldots,z_{m-1}),$ (1.6)
where $\oint_{0}$ denotes an integral along a circle around the origin with counterclockwise orientation and sufficiently small radius. The function $\mathbf{D}_{\mathfrak{h}}(z_{1},\ldots,z_{m-1})$ is defined as a Fredholm determinant in Definition 2.1. An equivalent definition through a series expansion will be discussed in Section 2.2.
@@
The function $\chi_{\mathfrak{h}}(\eta,\xi)$ is defined in terms of Brownian motion hitting expectations, an idea highly inspired by [MQR21]. Indeed, $\chi_{\mathfrak{h}}(\eta,\xi)$ should be understood as the Brownian hitting operators in [MQR21] written in Fourier-like spaces. Nevertheless, we stress that our results do not follow directly from [MQR21]. In the multi-time situation, direct connections to determinantal point processes and the Eynard-Mehta theorem are lost and the bi-orthoganalization procedure here arises differently and takes a different form. On the contrary, our results are, in some sense, more general. Indeed, if we set the time parameters to be the same (which is allowed in the assumption of the theorem), the right-hand side of (1.6) can be shown to recover the formulas in [MQR21], after some quite non-trivial manipulations. We refer to Section 6.2 for the details, see also [LO25] which treats the special narrow wedge case.
@@ 1.3 Outline of the proof and some discussions
A key ingredient of this paper is an explicit probabilistic representation of the function $\mathrm{ch}_{Y}(v,u)$, through a hitting expectation with respect to geometric random walks, see Theorem 3.4. The probabilistic representation is suitable for asymptotic analysis and leads to the Brownian hitting representation in the limit. For technical reasons, we first take the limit of the TASEP formula under the assumption that the KPZ fixed point starts with initial conditions consisting of finitely many narrow wedges, and then extend the formula to compactly supported initial condition at the level of the KPZ fixed point, using a density argument and the continuity of the law of the KPZ fixed point with respect to initial conditions.
Finally, we comment on our assumptions on the initial conditions. It would be desirable if one can get a formula that works for all initial conditions $\mathfrak{h}\in\mathrm{UC}$, in particular, the flat initial condition $\mathfrak{h}\equiv 0$. The reason we choose to restrict to the subspace $\mathrm{UC}_{c}$ is not merely a technical issue. There are genuine structural difficulties in this generality: the characteristic function $\chi_{\mathfrak{h}}(\eta,\xi)$ of the initial condition (see Definition 2.2) may not be well-defined pointwisely in general. Indeed, for the flat initial condition $\mathfrak{h}\equiv 0$, one can show from our formula that $\chi_{\mathfrak{h}}(\eta,\xi)$ is the dirac delta kernel $\delta_{\eta=-\xi}$, when properly interpreted. This suggests that in general, one should conjugate our formula to real spaces so that the limit when the support goes to infinity exists, at the level of operators acting on real spaces. We choose to stick to contour integral/Fourier type kernels, as it makes the algebraic structure of the multi-time formula much nicer and the analogy between narrow wedge and general initial conditions much more transparent. We leave it as a future project to extend our formula to any $\mathfrak{h}\in\mathrm{UC}$, with a kernel acting on real spaces.
@@
Throughout the paper, we will mostly use English letters $x,t,h,u,v,w,\ldots$ for the pre-limit (TASEP) formulas and Greek letters $\alpha,\tau,\beta,\xi,\eta,\zeta,\ldots$ for the limiting (KPZ fixed point) formulas. A detailed summary of the notation we use is in the following table.
@@
The rest of the paper is organized as follows. In Section 2 we present the formulas for the main part $\mathbf{D}_{\mathfrak{h}}$ appearing in Theorem 1.2, both as a Fredholm determinant in Section 2.1, and as a Fredholm series expansion in Section 2.2. Then in Section 3 we present and prove the corresponding pre-limit formulas for TASEP, in particular in Section 3.1 we prove that the characteristic function of the initial condition is given by a random walk hitting expectation. In Section 4 we prove convergence of the TASEP formulas to the KPZ fixed point formulas, under the assumption that the initial condition of the KPZ fixed point consists of finitely many narrow wedges. We then extend the KPZ fixed point formula to any compactly supported initial condition in Section 5. Finally, in Section 6 we show that at equal-time, our formula reduces to a genuine Fredholm determinant, which is then shown to be equivalent to the path integral formula of [MQR21].
@@ Acknowledgements
YL acknowledges the support of a Junior Fellowship at Institut Mittag-Leffler in Djursholm, Sweden, during the Fall semester of 2024, funded by the Swedish Research Council under grant no. 2021-06594. The work of ZL is partially supported by NSF DMS-2246683. The authors thank Konstantin Matetski, Li-Cheng Tsai, Xuan Wu, and Nikolaos Zygouras for their comments and suggestions.
@@ 2 Multipoint distribution formula for the KPZ fixed point
In this section, we explain in details the function $\mathbf{D}_{\mathfrak{h}}(z_{1},\ldots,z_{m-1})$ appearing on the right-hand side of (1.6). Proofs of the formula will be deferred to Section 4 and Section 5.
2.1 Fredholm determinant representation of $\mathbf{D}_{\mathfrak{h}}(z_{1},\ldots,z_{m-1})$
The function $\mathbf{D}_{\mathfrak{h}}$ is defined in the same way as its analog in [Liu22] for the narrow wedge initial condition. The only difference is the part involving the initial condition $\mathfrak{h}$, which only appears in the top-left corner of the integral kernel. Below, we introduce the Fredholm determinant representation of the function $\mathbf{D}_{\mathfrak{h}}$.
@@ 2.1 Fredholm determinant representation of $\mathbf{D}_{\mathfrak{h}}(z_{1},\ldots,z_{m-1})$
Let $\Gamma_{m,\mathrm{L}}^{\mathrm{out}}$, $\ldots$, $\Gamma_{2,\mathrm{L}}^{\mathrm{out}}$, $\Gamma_{1,\mathrm{L}}$, $\Gamma_{2,\mathrm{L}}^{\mathrm{in}}$, $\ldots$, $\Gamma_{m,\mathrm{L}}^{\mathrm{in}}$ be 2m12m-1 “nested” contours in the region $\mathbb{C}_{\mathrm{L}}$. They are all unbounded contours from $\infty\mathrm{e}^{-2\pi\mathrm{i}/3}$ to $\infty\mathrm{e}^{2\pi\mathrm{i}/3}$. Moreover, they are located from the right (corresponding to the superscript $\mathrm{out}$”) to the left ($\mathrm{in}$”). The superscripts $\mathrm{out}$” and $\mathrm{in}$” should be understood with respect to the point -$\infty$. Similarly, let $\Gamma_{m,\mathrm{R}}^{\mathrm{out}}$, $\ldots$, $\Gamma_{2,\mathrm{R}}^{\mathrm{out}}$, $\Gamma_{1,\mathrm{R}}$, $\Gamma_{2,\mathrm{R}}^{\mathrm{in}}$, $\ldots$, $\Gamma_{m,\mathrm{R}}^{\mathrm{in}}$ be 2m12m-1 “nested” contours from left to right on the half plane $\mathbb{C}_{\mathrm{R}}$. They are from $\infty\mathrm{e}^{-\pi\mathrm{i}/5}$ to $\infty\mathrm{e}^{\pi\mathrm{i}/5}$. Their superscripts $\mathrm{out}$” and $\mathrm{in}$” could be understood with respect to the point $\infty$. Note that the angles for the left contours and right contours are chosen differently. The choice of the angles guarantees super-exponential decay of the kernel along the contours even if $\tau_{i}=\tau_{i+1}$ for some $i$. See [Liu22, LZ25] for more discussions on the choices of the angles. See Figure 1 for an illustration of the contours.
@@
$\Gamma_{\ell,\mathrm{L}}:=\Gamma_{\ell,\mathrm{L}}^{\mathrm{out}}\cup\Gamma_{\ell,\mathrm{L}}^{\mathrm{in}},\qquad\Gamma_{\ell,\mathrm{R}}:=\Gamma_{\ell,\mathrm{R}}^{\mathrm{out}}\cup\Gamma_{\ell,\mathrm{R}}^{\mathrm{in}},\qquad\ell=2,\ldots,m,$
@@
${\mathrm{d}}\mu(\zeta)={\mathrm{d}}\mu_{\bm{z}}(\zeta):=\begin{dcases}\frac{-z_{\ell-1}}{1-z_{\ell-1}}\frac{{\mathrm{d}}\zeta}{2\pi{\mathrm{i}}},&\zeta\in\Gamma_{\ell,\mathrm{L}}^{\mathrm{out}}\cup\Gamma_{\ell,\mathrm{R}}^{\mathrm{out}},\quad\ell=2,\ldots,m,\\ \frac{1}{1-z_{\ell-1}}\frac{{\mathrm{d}}\zeta}{2\pi{\mathrm{i}}},&\zeta\in\Gamma_{\ell,\mathrm{L}}^{\mathrm{in}}\cup\Gamma_{\ell,\mathrm{R}}^{\mathrm{in}},\quad\ell=2,\ldots,m,\\ \frac{{\mathrm{d}}\zeta}{2\pi{\mathrm{i}}},&\zeta\in\Gamma_{1,\mathrm{L}}\cup\Gamma_{1,\mathrm{R}}.\end{dcases}$ (2.2)
@@
$\mathrm{D}_{\mathfrak{h}}(z_{1},\ldots,z_{m-1})=\det\left(\mathrm{I}-\mathrm{K}_{1}\mathrm{K}_{\mathfrak{h}}\right),$
@@
$\mathrm{f}_{i}(\zeta):=\begin{dcases}\mathrm{e}^{-\frac{1}{3}(\tau_{i}-\tau_{i-1})\zeta^{3}+(\alpha_{i}-\alpha_{i-1})\zeta^{2}+(\beta_{i}-\beta_{i-1})\zeta},&i=2,\ldots,m,\\ \mathrm{e}^{-\frac{1}{3}\tau_{1}\zeta^{3}+\alpha_{1}\zeta^{2}+\beta_{1}\zeta},&i=1.\end{dcases}$ (2.7)